alladega Tire and Rubber Company has capacity to produce 119,000 tires. Talladega presently produces and sells 91,000 tires for the North American market at a price of $102 per tire. Talladega is evaluating a special order from a European automobile company, Autobahn Motors. Autobahn is offering to buy 14,000 tires for $84.4 per tire. Talladega’s accounting system indicates that the total cost per tire is as follows:

Direct materials $39
Direct labor 14
Factory overhead (60% variable) 23
Selling and administrative expenses (40% variable) 20
Total $96

Talladega pays a selling commission equal to 5% of the selling price on North American orders, which is included in the variable portion of the selling and administrative expenses. However, this special order would not have a sales commission. If the order was accepted, the tires would be shipped overseas for an additional shipping cost of $6 per tire. In addition, Autobahn has made the order conditional on receiving European safety certification. Talladega estimates that this certification would cost $75,600.

a. Prepare a differential analysis dated July 31 on whether to reject (Alternative 1) or accept (Alternative 2) the special order from Autobahn Motors. If an amount is zero, enter "0". If required, round interim calculations to two decimal places.

Differential Analysis
Reject Order (Alt. 1) or Accept Order (Alt. 2)
July 31
Reject Order (Alternative 1) Accept Order (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $ $ $
Costs:
Direct materials
Direct labor
Variable factory overhead
Variable selling and admin. expenses
Shipping costs
Certification costs
Income (Loss) $ $ $

b. What is the minimum price per unit that would be financially acceptable to Talladega? Round your answer to two decimal places.
$per unit
Answer #1

a.

Differential Analysis
Reject Order (Alt. 1) or Accept Order (Alt. 2)
July 31
Reject Order (Alternative 1) Accept Order (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $0 14000*84.40 = $1181600 $1181600
Costs:
Direct materials 0 14000*39 = $546000 (546000)
Direct labor 0 14000*14 = $196000 (196000)
Variable factory overhead 0 14000*23*60% = $193200 (193200)
Variable selling and admin. expenses 0 14000*20*40% - 102*14000*5% = $40600 (40600)
Shipping costs 0 14000*6 = $84000 (84000)
Certification costs 0 75600 (75600)
Income (Loss) $0 $46200 $46200

b. Minimum price per unit = $84.40 - 46200/14000 = $81.10 per unit